A closed box, whose length in twice its width, is to have a surface area of 192 square inches. Find the dimensions of the box when the volume is maximum?


1
Expert's answer
2021-01-11T18:03:58-0500

Let "l,w,h" be the dimensions of the closed box.

Given: "l=2w" and Surface area "=192" square inches.

"\Rightarrow 2lw+2wh+2lh=192"

"\Rightarrow lw+wh+lh=96"

"\Rightarrow2w^2+wh+2wh=96"

"\Rightarrow2w^2+3wh=96"

"\Rightarrow h=\frac{96-2w^2}{3w}" ... (1)

Now volume of the box is "V=lwh"

"\Rightarrow V=(lw)\frac{96-2w^2}{3w}" ( from (1) )

"\Rightarrow V=\frac{96l-2lw^2}{3}"

Let "f(l,w)=\frac{96l-2lw^2}{3}"

Find the first order partial derivatives of the above function.

"f(l,w)=\frac{96l-2lw^2}{3}\Rightarrow f_{l}=\frac{96-2w^2}{3}" and "f_{w}=\frac{0-4lw}{3}"

"\Rightarrow f_{l}=\frac{96-2w^2}{3}" and "f_{w}=\frac{-4lw}{3}"

For maximum volume, set "f_{l}=0" and "f_{w}=0"

"\Rightarrow \frac{96-2w^2}{3}=0" and "\frac{-4lw}{3}=0"

"\Rightarrow 2w^2=96" and "l=0" or "w=0"

"\Rightarrow w^2=48" ( since "l\neq 0,w\neq 0" )

"\Rightarrow w=4\sqrt{3}"

Substituting "w=4\sqrt{3}" in equation (1), we get "h=0"

And "w=4\sqrt{3}\Rightarrow l=8\sqrt{3}"

Therefore, dimensions of the closed box are: "l=8\sqrt{3}" inches and "w=4{\sqrt{3}}" inches